# Traffic Response Functions: Patterns, Propagation and Congestion

### v2

@@ Abstract
We further pursue a recently tested approach to the study of transport systems. Using empirical data collected on German motorways, we investigate the use of response functions as a tool to study traffic dynamics in motorway networks. We use a response function that measures the change in traffic observed at a given location over time conditioned on congestion at another location. By introducing several velocity-dependent congestion indicators, we are able to work out the basic characteristics of the flow and density responses to congestion. Our results demonstrate that these responses capture the propagation of congestion, the acceleration phases of traffic participants during congestion, and the correlation between the three fundamental traffic observables: flow, density, and velocity. Furthermore, a direct comparison with the underlying data reveals the potential for characterizing traffic patterns with response functions. A relationship between the shape of the response functions and the traffic oscillation during a traffic jam can be demonstrated. Moreover, our findings allow us to determine the propagation velocity of congestion from responses. Importantly, in the course of our analysis, we also identify a scaling property that results in the collapse of the response functions onto a single curve.
@@ 1 Introduction
Vehicular traffic on a multi-lane motorway is a highly complex and dynamic process. Millions of drivers experience this complexity on a daily basis due to a rather simple and seemingly unavoidable fact: traffic congestion. Congestion is a phenomenon arising in a collectively driven many-body system that unfolds its complex dynamics in different scales of space and time. Over the past six decades, a variety of theories and models have been developed to comprehend and describe the complex, non-linear behavior of traffic systems. These include fluid-dynamical theories [1, 3], gas-kinematic models [2, 5], car-following theories [4, 6, 11], cellular automata models [7, 8, 9, 10], as well as the three-phase traffic theory [12, 13, 14], and others. In view of the central role played by vehicular traffic in the mobility and transportation of modern societies, alternative research methodologies may facilitate a more comprehensive understanding of these systems and their dynamics, thereby enabling the development of new tools for effective traffic planning, management and control.
We have recently conducted a number of studies along these lines, focusing on the application of concepts and approaches used to study financial markets in the context of traffic networks. We conducted a comprehensive analysis of traffic data to investigate temporal [15] and spatial [17] correlations, leading to the identification of dominant and subdominant collectivities [16, 18] and new characteristics of traffic networks. The most recent work [19] explores the potential of response functions as a tool for studying traffic dynamics. The fundamental premise of the concept of response functions is the ability to examine a system’s reaction to specific events, thereby revealing the influence of these events on the response variable and the sensitivity of the system to such influences.
In the field of finance, response functions have been successfully applied to analyse the reaction of a system to specific events on different time scales, highlighting the complexity of market behaviors. Previous trades and correlated sector responses drive present and future price dynamics. In this context, response functions have enabled significant advances in our understanding of cross-impact mechanisms [21, 23, 24], liquidity effects [23, 25], sectoral dependencies [22], and the non-Markovian nature of price fluctuations [20] in financial markets. In the context of transportation networks, congestion is one of the most significant events to study the responses of traffic observables. By selecting velocity as the response variable, response functions can identify the presence of congestion, its propagation along motorway segments, and therefore spatiotemporal characteristics of traffic patterns [19]. These findings establish a crucial foundation for studying traffic networks with the concept of response functions. Full characterization of traffic patterns is a challenging task, and information on other fundamental observables, such as density and flow, is valuable.
The aim of this contribution is to further establish response functions as a research instrument for the study of traffic dynamics. To this end, we extend the previous approach presented by Wang et al. [19] by also incorporating flow and density responses for spatially subsequent motorway sections. By comparing all three types of responses, our objective is to identify the fundamental characteristics of flow and density responses. Additionally, we define multiple congestion indicators according to different velocity ranges. The aim is to identify different phases or patterns of congestion. In this process, we also investigate the potential for deriving the propagation velocity of congestion using responses.
@@ 5 Conclusion
The response functions of the fundamental traffic observables velocity, flow and density were analyzed based on congestion indicators for different velocity ranges and spatial subsequent sections on heavily trafficked motorways in NRW, Germany. All three response types of observables exhibited clear local minima (maxima) or valleys (hills) within the first 30 minutes of time lags, thereby demonstrating the impact of congestion and its propagation over several kilometers. The correlation between the three fundamental observables is reflected in the responses. Furthermore, the importance of considering multiple indicators with different velocity ranges to indicate congestion was demonstrated. Additionally, the propagation of congestion through on- and off-ramps of interchanges into other parts of the motorway network was verified.
Moreover, the responses capture the acceleration-deceleration cycles during congestion, which reflects traffic oscillations, or stop-and-go patterns. In view of this finding, we proceeded to differentiate between congestion patterns with the aid of response functions. A sample of the data was considered, and it was found that the width of the minima (maxima) or valleys (hills) in the course of the responses coincides with the average period of the occurring traffic oscillations. Furthermore, we narrowed down congestion phases according to three-phase traffic theory, with information gathered from the responses. Additionally, we derived the propagation velocity of congestion from the responses. Despite the numerical results aligning well with empirical observations, it was not feasible to derive a universal relationship between the propagation velocities and the distance to the indicator section. Lastly, motivated by the similarities in the shape of the responses for small time shifts, we successfully conducted an approach for demonstrating a scaling behaviour of the responses.
Our study faced several limitations, primarily related to data quality, detector availability, and spatial constraints. Data quality significantly influenced the observed characteristics and variations, with one-minute intervals in detector measurements limiting our analysis. In addition, differentiating congestion patterns proved challenging; the averaged data limited the ability to analyze the relationship between the widths of minima and traffic oscillations. The determination of propagation velocities and scaling behavior is highly dependent on the distinct manifestations of local minima. These limitations underscore the locally constrained nature of our approach and suggest that improvements in data quality and detector distribution may enhance future analyses.
Despite the limitations, the results of our paper demonstrate that response functions are useful tools to study local traffic dynamics and to analyze emerging traffic patterns. Responses reveal the change in traffic observables at particular road sections based on previous congestion at other sections. Such information is very helpful in traffic engineering applications such as estimating, tracking or predicting traffic patterns. A major advantage of the response approach is that it avoids the complex calculations and parameter predictions required to apply different models. In addition, the presented approach can be carried out with a higher level of detail. It is likely that a more accurate analysis of traffic patterns can be made by determining the responses for each individual lane. This is of particular interest with regard to the formation of congestion patterns, their characterisation and the determination of their propagation velocity. Possible follow-up studies in this direction seem promising to uncover further potential of response functions as a useful tool for traffic scientists and engineers.

### v3

# Correlations in Motion: A Simple Response-Based Analysis of Traffic Flow

@@ Abstract
Why does a traffic jams form out of nowhere, and why does it stretch for kilometers even after the initial cause is passed? This study examines how congestion moves and spreads across motorways using a surprisingly simple method: response functions. These functions are based purely on data and show how changes in traffic flow, density, and velocity are connected over time and space. Using real-world data from German motorways, we track how traffic reacts to earlier disturbances, capturing the waves of slowing and accelerating that drivers experience in stop-and-go traffic. The results demonstrate how congestion propagates and how its rhythm can be measured and predicted. Unlike complex traffic models, this approach requires no simulations or assumptions about driver behavior. It works directly from the information provided by the road. The goal is clear: to understand congestion better so that we can manage it more effectively and perhaps spend less time stuck in it.
@@ 1 Introduction
Traffic congestion is a widespread feature of modern mobility that affects the daily routines of millions. Yet, its emergence and evolution remain difficult to fully predict and control. On motorways in particular, traffic behaves like a complex, many-body system where local disruptions can lead to widespread congestion. Despite advances in traffic science, understanding how disturbances travel across a network and how they affect key traffic variables, such as flow, velocity, and density, remains a major challenge.
Everyone who has ever been caught in a traffic jam has likely asked the same questions: Where is this coming from? Why does it spread? And why does it sometimes dissolve just as quickly as it formed? The experience of driving can be surprisingly disconnected from the bigger picture of what is actually happening along the road network. A slowdown in one area can spread through the system, affecting traffic kilometers away. However, understanding how and why this happens, and how one traffic variable influences another, is not trivial. Scientists have long known that variables such as velocity, vehicle density, and flow are closely linked. Yet, knowing that a link exists is not the same as being able to measure it clearly.
This raises a fundamental question: How can we quantify these interdependencies and understand how congestion propagates in time and space, without relying on overly complex models? Can we capture these dynamics using only data, and do so in a way that is robust, transparent, and computationally efficient? A method that meets these criteria could benefit not only researchers, but also planners and engineers, by providing practical tools for interpreting traffic data and identifying emerging patterns. It would also support the broader effort to move from descriptive observations to measurable and actionable findings in traffic systems research.
To address this challenge, we turn to response functions, a method that measures how changes in traffic observables at one location follow congestion events at another. Applied in financial market analysis, response functions quantify the impact of past events on a system’s behavior across different time scales. In finance, response functions have been used to reveal the complexity of market reactions and the temporal structure of cross-asset dependencies. Notably, they have enabled significant advances in the understanding of cross-impact mechanisms [2, 4, 5], liquidity effects [4, 6], sectoral dependencies [3], and the non-Markovian nature of price fluctuations [1]. Previous trades and correlated sector responses influence future price movements in measurable and often nonlinear ways. These findings motivate the use of similar techniques in other complex systems, including transportation.
By conditioning responses on defined congestion indicators, we can extract how congestion affects downstream traffic, and do so in a way that reveals both propagation velocity and oscillation patterns, all from empirical time series data. While many established traffic models, such as fluid-dynamical theories [7, 9], gas-kinetic approaches [8, 11], and car-following theories [10, 12, 17], offer understanding of the mechanics of traffic flow, they often involve complex parameter estimation and limited generalizability. Similarly, cellular automata [13, 14, 15, 16] and three-phase traffic theory [18, 19, 20] have successfully reproduced emergent congestion patterns, but at the cost of considerable model complexity.
In contrast, response functions offer a complementary perspective: a straightforward way to empirically quantify how traffic reacts to prior disruptions. Recent studies have shown their potential in capturing temporal and spatial effects of congestion using velocity as the primary observable [21, 23, 25]. However, their broader capacity to quantify the interplay between flow, density, and velocity, and to systematically assess the spatiotemporal spread of congestion, remains largely unexplored. In this study, we address that by applying response functions to real motorway data from North Rhine–Westphalia, Germany, and investigate their ability to characterize traffic dynamics across multiple observables.
@@ 5 Conclusion
Traffic congestion affects millions of drivers daily, causing delays, frustration, and lost time. While this study focuses on the complex dynamics underlying congestion, the findings gained have direct practical value for every traffic participant. By quantifying how congestion propagates and how key traffic variables influence one another, we provide tools that can help improve traffic management, prediction, and control, ultimately aiming to reduce the time drivers spend stuck in traffic congestion.
Our analysis of response functions based on empirical data from heavily trafficked motorways in North Rhine-Westphalia revealed clear patterns of congestion propagation across several kilometers. The close interplay between velocity, flow, and density emerged naturally from the data, and response functions capture the oscillatory stop-and-go behavior familiar to many drivers. These response features showed distinct local minima and maxima within the first 30 minutes after congestion events, demonstrating the immediate impact and temporal evolution of traffic disturbances. By examining the width and shape of these features, we identified characteristic time scales that correspond to traffic oscillations and are consistent with phases described by three-phase traffic theory. This allowed us to differentiate congestion states and better understand their dynamics. Additionally, we derived estimates for the propagation velocity of congestion waves traveling through the motorway network. While these velocities align well with empirical observations, a universal relationship between propagation velocity and the distance from the initial congestion could not be identified. Another important finding is the discovery of a scaling property in the response functions, which causes a collapse of multiple response curves onto a single universal pattern. This hints at underlying regularities in traffic dynamics.
Several limitations influenced our study, primarily related to data quality, detector availability, and spatial constraints. The one-minute resolution of detector measurements limited the granularity of the analysis, and the spatial distribution of detectors restricted the ability to track congestion over larger areas. Moreover, averaging over time and lanes reduced the clarity of certain traffic patterns, making it challenging to precisely differentiate between congestion types or oscillation details. These factors affected the sharpness of local minima and maxima in the responses, which are critical for determining propagation velocities and scaling behavior. Despite these constraints, the response function approach offers a robust, data-driven alternative to more complex, parameter-heavy traffic models.
Importantly, the method holds significant promise for practical applications. Response functions provide a relatively simple yet powerful way to characterize local traffic dynamics and emerging patterns in real time. They reveal how traffic observables at specific road sections change based on prior congestion elsewhere, enabling better estimation, tracking, and potentially prediction of traffic conditions. Unlike traditional models, this approach does not require extensive calibration or assumptions about driver behavior, making it more adaptable and transparent. Further improvements, such as analyzing responses on a lane-by-lane basis or integrating more comprehensive detector networks, could increase accuracy and the ability to resolve detailed congestion phenomena. Such advances would directly support traffic engineers and planners in designing more effective traffic management strategies.
@@ Acknowledgements
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used ChatGPT in order to improve the language and readability. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.